# Efficient Interleaved Speech Modeling through Knowledge Distillation

### v1

@@ Abstract
Efficient Interleaved Speech Modeling through Knowledge Distillation
Mohammadmahdi Nouriborji Nlpie Research m.nouriborji@nlpie.com                        Morteza Rohanian University of Zurich morteza.rohanian@uzh.ch
@@ 1 Introduction
A more tractable but limited alternative assembles an automatic speech recognition (ASR) frontend, a text-only LLM, and a text-to-speech (TTS) backend. While effective for basic language tasks, this pipeline discards much of the expressive signal—prosody, speaker identity, and acoustic nuance—leading to unnatural outputs. Recent speech models address this by modeling audio end-to-end using discrete units. However, most of these models are trained from scratch at large scales, and model providers often maintain separate checkpoints for each deployment size. This practice is resource-intensive and inefficient, especially when targeting multiple operating points.
Figure 1: Distillation framework: The student aligns with the teacher’s logits, intermediate states, and ground-truth labels, guided by a multi-part loss.
We introduce TinyWave, a family of 2 billion-parameter speech-to-speech transformers distilled from the larger SpiritLM Nguyen et al. (2025). TinyWave models preserve expressive qualities of speech and support both audio-only and interleaved speech–text inputs. Instead of training new models from scratch, we use layer-aligned distillation to transfer knowledge from a high-capacity teacher. We match hidden states, attention maps, and softened logits between teacher and student, encouraging the student to learn the teacher’s internal mechanisms. We also fine-tune the teacher on target-domain data to reduce distribution mismatch before distillation.
@@
All experiments use the publicly available Libri-Light dataset, which represents only a fraction of the data used to train the teacher model Nguyen et al. (2025). Our results show that high-quality expressive speech modeling is possible without access to private datasets or large-scale compute.
In summary, our work contributes:
@@
    Open-source training and evaluation tools that enable reproducibility, extension, and deployment on commodity hardware.
@@ 2 Methodology
We present the TinyWave family: compact 2-billion-parameter speech-to-speech models distilled from 7B-parameter SPIRIT LM teachers Nguyen et al. (2025). The SPIRIT LM models extend LLaMA-2 with speech capabilities by training on continuous streams of HuBERT-based audio tokens and interleaved text. TinyWave models inherit this cross-modal foundation through distillation. We release two speech-only variants and one interleaved model that handles mixed audio-text sequences. Our training pipeline combines model initialization, teacher correction, and a layer-to-layer distillation framework.
@@ 2.2 Speech-Only Models
For the base speech variant, we fine-tune the SpiritLM-Base-7B teacher on a 20% subset of Libri-Light. This teacher correction step ensures the supervision matches the target domain and avoids distributional mismatch. We skip this step for the expressive-speech model, since its teacher is already trained to encode prosodic and expressive cues relevant to the target task.
@@ 2.3 The Interleaved Model
We sample from five interleaving patterns during training: [Speech], [Text], [Speech][Text], [Text][Speech], and [Speech][Text][Speech], each with a fixed probability. This sampling encourages diversity and teaches the model to switch between modalities mid-sequence. We fine-tune the 7B expressive teacher on this interleaved corpus for a short duration (10k steps), then distill its knowledge into the student using our distillation method.
@@ 2.4 Layer-to-Layer Distillation
$\mathcal{L}_{\text{output}}=\text{KL}\left(\text{softmax}\left(\frac{y^{(t)}}{\tau}\right)\|\text{softmax}\left(\frac{y^{(s)}}{\tau}\right)\right)$ (2)
Finally, we apply a cross-entropy loss over ground-truth tokens:
$\mathcal{L}_{\text{LM}}=-\sum_{t}\log P(y_{t}\mid y_{<t})$ (3)
The full objective combines all three components:
The hyperparameters $\lambda_{1}$, $\lambda_{2}$, and $\lambda_{3}$ were tuned via grid search over [0.1, 0.5, 1.0], with $\tau=2.0$ for softened logits, balancing alignment, output matching, and language modeling objectives. This loss enables the student to replicate the teacher’s behavior efficiently, making TinyWave suitable for deployment in constrained environments without compromising expressive generation capabilities.
Table 1: Normalised Perplexity Score (↑).
@@ 3 Results
We evaluate TinyWave across multiple dimensions: (i) language modeling quality (NPS), (ii) spoken commonsense reasoning (StoryCloze), (iii) morphosyntactic competence (sWUGGY, sBLIMP), (iv) acoustic consistency and alignment (SALMon), and (v) task-level transfer through joint ASR and TTS training. An ablation study quantifies the impact of individual distillation components. Throughout, Teacher refers to the original 7 B SpiritLM checkpoint, Distilled to the 2 B TinyWave student, and Baseline to an equally sized model trained from scratch without distillation. In addition to the teacher model, we compare our approach with several strong baselines including GSLM Lakhotia et al. (2021), AudioLM Borsos et al. (2023), VoxtLM Maiti et al. (2024), TWIST Hassid et al. (2023), and LAST-1.3 Turetzky and Adi (2025).
Language-Modeling Quality
Spoken Commonsense Reasoning
Surface-Level Morphosyntax
Acoustic Consistency and Alignment
Table 3: SALMon accuracy (↑).
@@
Multimodal Feature Transfer with ASR and TTS
To evaluate TinyWave’s capacity for real-world speech tasks, we jointly fine-tune the interleaved model on automatic speech recognition (ASR) and text-to-speech (TTS) using Low-Rank Adaptation (LoRA). We integrate LoRA modules into both the attention and feedforward (MLP) sublayers of the TinyWave architecture, enabling it to specialize in both speech decoding and synthesis simultaneously. Despite the model’s compressed size, the fine-tuned TinyWave-Interleaved retains strong generalization: it achieves a word error rate (WER) of 0.334 on ASR and a character error rate (CER) of 0.142 on TTS (Table 4). This corresponds to an increase of 3.3 absolute WER points and 3.9 CER points over the 7B-parameter teacher model.
Distillation Components
To understand the importance of each component in our distillation strategy, we conduct an ablation study isolating the contributions of output-level and internal supervision. Table 5 compares three configurations: (i) a version inspired by Minitron muralidharan2024compact that uses only softened output logits; (ii) a DistilBERT-style setup with logits-based supervision but no intermediate alignment; and (iii) our full framework that combines both output and layer-to-layer supervision. The results show that removing layer-to-layer alignment leads to a 2–3 point drop in NPS. When relying solely on logits, the gap widens further—exceeding 4 points. Our method achieves the highest NPS, confirming that deep supervision across layers significantly enhances the quality and fidelity of the distilled model.
@@ 4 Related Works
LLMs have recently been extended to operate on waveform tokens. This enables high-quality multi-speaker speech synthesis Wang et al. (2023a); Kharitonov et al. (2023); Jiang et al. (2024); Ji et al. (2024) and general audio generation Kreuk et al. (2023); Borsos et al. (2023). Models like VALL-E and AudioLM separate phonetic and acoustic information to improve fidelity Wang et al. (2023a); Borsos et al. (2023).
Self-supervised speech models—such as wav2vec 2.0, HuBERT, and w2v-BERT—can be discretized into pseudo-phonemes. These discrete tokens allow LMs to learn over raw audio Lakhotia et al. (2021). The tokens capture both linguistic content and prosody Kharitonov et al. (2022), supporting tasks like emotion conversion Kreuk et al. (2023) and spoken dialogue generation Nguyen et al. (2023). However, under equal data budgets, speech-only LMs still lag text LMs in semantic accuracy Nguyen et al. (2023).
GSLM Lakhotia et al. (2021) marked the start of large-scale generative speech models. Successors such as Spectron Nachmani et al. (2024), AudioPaLM Rubenstein et al. (2023), VioLA Wang et al. (2023b), VoxtLM Maiti et al. (2024), and SUTLM Chou et al. (2023) unify ASR, TTS, and spoken continuation. Despite impressive capability, these models exceed 7B parameters and are compute-heavy.
Knowledge distillation addresses this bottleneck. Originally, it involved matching softened output logits Hinton et al. (2015). Later approaches aligned hidden states and attention maps Sanh et al. (2019); Sreenivas et al. (2024). In some cases, student models outperform their teachers due to improved regularization.
In speech, most distillation efforts have focused on ASR encoders Gandhi et al. (2023). End-to-end speech generation distillation is largely unexplored—this is the gap TinyWave fills.
TinyWave compresses large audio LLMs into a 2B student using a layer-wise distillation scheme. It retains near-teacher quality while enabling efficient inference on commodity hardware.
Table 4: ASR (WER ↓) and TTS (CER ↓).
Model ASR WER TTS CER
Teacher (7 B) 0.301 0.103
TinyWave-Interleaved 0.334 0.142
Table 5: Ablation on NPS (↑).
Setup Logits Layer-to-Layer NPS
Minitron 0.826
DistilBERT 0.840
Ours 0.865
@@ 5 Limitations
Despite the promising performance of TinyWave models across semantic and expressive benchmarks, several limitations remain. Notably, the models exhibit constrained semantic understanding, often prioritizing phonetic accuracy over deeper linguistic comprehension. This limitation is consistent with findings in self-supervised speech models, which tend to encode phonetic information more robustly than semantic content. Incorporating more advanced language models, such as the latest iterations of LLaMA, could enhance the semantic capabilities of speech generation systems.
The current tokenizer and vocoder components present challenges in producing natural, human-like speech. Existing tokenizers may inadequately capture the stylistic nuances of conversation, leading to outputs that lack expressiveness and emotional depth. Vocoders often struggle with prosody—the rhythm, stress, and intonation of speech—resulting in synthetic speech that sounds flat or unnatural. Future work should explore the integration of advanced tokenization methods, such as Residual Vector Quantization (RVQ), and more sophisticated vocoders capable of modeling fine-grained prosodic features to improve the naturalness and expressiveness of generated speech.
@@ 6 Conclusion
TinyWave compresses a 7 B speech–text foundation model into 2 B parameters without discarding expressive power. Layer-aligned distillation, anchored by a brief teacher-correction phase, transfers both semantic reasoning and prosodic control to the student. Across NPS, StoryCloze, SALMon, and ASR/TTS, TinyWave matches or closely trails its teacher and consistently outperforms size-matched baselines. These results show that careful distillation, not additional scale, can deliver practical, high-quality speech generation on modest hardware. We release all checkpoints and training code to support reproducibility and future work.

### v2

@@ Abstract
Index Terms—  speech generation, speech language models, knowledge distillation, multimodal transformers
@@ 1 Introduction
A more tractable but limited alternative assembles an automatic speech recognition (ASR) frontend, a text-only LLM, and a text-to-speech (TTS) backend. While effective for basic language tasks, this pipeline discards much of the expressive signal leading to unnatural outputs. Recent speech models address this by modeling audio end-to-end using discrete units. Most of these models are trained from scratch at large scales, and model providers often maintain separate checkpoints for each deployment size. This practice is resource-intensive and inefficient, especially when targeting multiple operating points.
Fig. 1: Distillation framework: The student aligns with the teacher’s logits, intermediate states, and ground-truth labels, guided by a multi-part loss.
We introduce TinyWave, a family of 2 billion-parameter speech-to-speech transformers distilled from the larger SpiritLM [14]. TinyWave models preserve expressive qualities of speech and support both audio-only and interleaved speech–text inputs. Instead of training new models from scratch, we use layer-aligned distillation to transfer knowledge from a high-capacity teacher. We match hidden states, attention maps, and softened logits between teacher and student, encouraging the student to learn the teacher’s internal mechanisms. We also fine-tune the teacher on target-domain data to reduce distribution mismatch before distillation.
@@
All experiments use the publicly available Libri-Light dataset, which represents only a fraction of the data used to train the teacher model [14]. Our results show that high-quality expressive speech modeling is possible without access to private datasets or large-scale compute.
In summary, our work makes the following contributions:
@@
    Open-source training and evaluation tools that enable reproducibility and deployment on commodity hardware.
@@ 2 Methodology
We present the TinyWave family: compact 2-billion-parameter speech-to-speech models distilled from 7B-parameter SPIRIT LM teachers [14]. The SPIRIT LM models extend LLaMA-2 with speech capabilities by training on continuous streams of HuBERT-based audio tokens and interleaved text. TinyWave models inherit this cross-modal foundation through distillation. We release two speech-only variants and one interleaved model that handles mixed audio-text sequences.
@@ 2.2 Speech-Only Models
For the base speech variant, we fine-tune the SpiritLM-Base-7B teacher on a 20% subset of Libri-Light. This teacher correction step ensures the supervision matches the target domain and avoids distributional mismatch. We skip this step for the expressive-speech model, since its teacher is already trained to encode prosodic and expressive cues relevant to the target.
@@ 2.3 The Interleaved Model
We sample from five interleaving patterns: [Speech], [Text], [Speech][Text], [Text][Speech], and [Speech][Text][Speech], each with a fixed probability. This sampling encourages diversity and teaches the model to switch between modalities mid-sequence. We fine-tune the 7B expressive teacher on this interleaved corpus for a short duration (10k steps), then distill its knowledge into the student using our distillation method.
@@ 2.4 Layer-to-Layer Distillation
Table 1: Normalised Perplexity Score (↑).
$\mathcal{L}_{\text{output}}=\text{KL}\left(\text{softmax}\left(\frac{y^{(t)}}{\tau}\right)\hskip 2.5pt\|\hskip 5.0pt\text{softmax}\left(\frac{y^{(s)}}{\tau}\right)\right)$ (2)
Finally, we apply a cross-entropy loss over ground-truth tokens:
$\mathcal{L}_{\text{LM}}=-\sum_{t}\log P(y_{t}\mid y_{<t})$ (3)
The full objective combines all three components:
This loss enables the student to replicate the teacher’s behavior efficiently, making TinyWave suitable for deployment in constrained environments without compromising expressive generation capabilities.
@@ 3 Results
We evaluate TinyWave across multiple dimensions: (i) language modeling quality (NPS), (ii) spoken commonsense reasoning (StoryCloze), (iii) morphosyntactic competence (sWUGGY, sBLIMP), and (iv) acoustic consistency and alignment (SALMon) [22]. Throughout, Teacher refers to the original 7 B SpiritLM checkpoint, Distilled to the 2 B TinyWave student, and Baseline to an equally sized model trained from scratch without distillation. In addition to the teacher model, we compare our approach with several strong baselines including GSLM [11], AudioLM [1], VoxtLM [12], TWIST [4], and LAST-1.3 [19].
@@
Table 3: SALMon accuracy (↑).
3.1 Language-Modeling Quality
3.2 Spoken Commonsense Reasoning
3.3 Surface-Level Morphosyntax
3.4 Acoustic Consistency and Alignment
@@ 4 Related Works
LLMs have recently been extended to operate on waveform tokens. This enables high-quality multi-speaker speech synthesis [20, 8, 7, 6] and general audio generation [10, 1]. Models like VALL-E and AudioLM separate phonetic and acoustic information to improve fidelity [20, 1].
Self-supervised speech models—such as wav2vec 2.0, HuBERT, and w2v-BERT—can be discretized into pseudo-phonemes. These discrete tokens allow LMs to learn over raw audio [11]. The tokens capture both linguistic content and prosody [9], supporting tasks like emotion conversion [10] and spoken dialogue generation [15]. However, under equal data budgets, speech-only LMs still lag text LMs in semantic accuracy [15].
GSLM [11] marked the start of large-scale generative speech models. Successors such as Spectron [13], AudioPaLM [16], VioLA [21], VoxtLM [12], and SUTLM [2] unify ASR, TTS, and spoken continuation. Despite impressive capability, these models exceed 7B parameters and are compute-heavy.
Knowledge distillation addresses this bottleneck. Originally, it involved matching softened output logits [5]. Later approaches aligned hidden states and attention maps [17, 18]. In some cases, student models outperform their teachers due to improved regularization. In speech, most distillation efforts have focused on ASR encoders [3]. TinyWave compresses large audio LLMs into a 2B student using a layer-wise distillation scheme. It retains near-teacher quality while enabling efficient inference on commodity hardware.
@@ 5 Limitations
Despite the promising performance of TinyWave models across semantic and expressive benchmarks, several limitations remain. Notably, the models exhibit constrained semantic understanding, often prioritizing phonetic accuracy over deeper linguistic comprehension. This limitation is consistent with findings in self-supervised speech models, which encode phonetic information more robustly than semantic content.
The current tokenizer and vocoder components present challenges in producing natural, human-like speech. Existing tokenizers may inadequately capture the stylistic nuances of conversation, leading to outputs that lack expressiveness and emotional depth. Vocoders often struggle with prosody resulting in synthetic speech that sounds flat or unnatural. Future work should explore the integration of advanced tokenization methods, such as Residual Vector Quantization (RVQ), and more sophisticated vocoders capable of modeling fine-grained prosodic features to improve the naturalness and expressiveness of generated speech.
@@ 6 Conclusion
TinyWave compresses a 7 B speech–text foundation model into 2 B parameters without discarding expressive power. Layer-aligned distillation, anchored by a brief teacher correction phase, transfers both semantic reasoning and prosodic control to the student. Across NPS, StoryCloze, SALMon, and ASR/TTS, TinyWave matches or closely trails its teacher and consistently outperforms size-matched baselines. These results show that careful distillation, not additional scale, can deliver practical, high-quality speech generation on modest hardware. We release all checkpoints and training code to support reproducibility and future work.
7 Acknowledgements
No funding was received for conducting this study. The authors have no relevant financial or nonfinancial interests to disclose.
8 Compliance with Ethical Standards
This is a numerical simulation study for which no ethical approval was required.